# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ Appendix I Computational Speed-Ups
The experiments were conducted on NVIDIA RTX A6000 GPUs with 48GB of memory. The complete synthetic experiment, spanning four datasets, took approximately 30 minutes to run. For the real survival data, each data fold required about 20 minutes, while each fold in the ablation study took around 40 minutes. All folds and datasets were run in parallel across multiple GPUs.
Code availability.
@@
Traditional survival models often impose parametric or semi-parametric assumptions on the hazard function. The Accelerated Failure Time (AFT) model [6] assumes a linear relationship between covariates and the logarithm of survival time, with parametric baseline distributions (e.g., Weibull). While interpretable, such models struggle with high-dimensional data and nonlinear covariate effects. The Cox Proportional Hazards (CoxPH) model [7], a semi-parametric approach, avoids specifying the baseline hazard but assumes proportional hazards. Though widely adopted, CoxPH’s linear predictor and proportionality constraints limit its flexibility in complex data regimes.
@@
To mitigate parametric assumptions, non-parametric methods like Random Survival Forests (RSF) [19] and Survival Support Vector Machines (SSVM) [36] emerged. RSF leverages ensemble learning for risk stratification but faces challenges in high-dimensional settings due to greedy tree induction. GP survival models [10] offer flexibility by modeling the hazard function nonparametrically, with inherent uncertainty quantification. Existing work has sought to address the cubic complexity in sample size of GPs by introducing variational inference techniques [21]. However, GPs remain fundamentally limited in scalability, particularly struggling with high-dimensional inputs and lacking the capacity to learn hierarchical representations, such as those required in image-based tasks [38].
@@
The advent of deep learning revolutionized survival analysis by enabling automatic feature learning from high-dimensional inputs. DeepSurv [20] extended CoxPH with neural networks, while DeepHit [28] employed multi-task learning for competing risks via discrete-time hazards. Discrete-time methods, including MTLR [43] and PCHazard [24], discretize the time axis to simplify likelihood computation, with recent advances like DySurv [32] incorporating conditional variational inference for dynamic prediction. Despite their predictive prowess, these models rely on frequentist training, yielding point estimates without uncertainty quantification — a significant shortcoming in safety-critical applications. Comprehensive reviews [42] highlight the rapid growth of deep survival methods but underscore their neglect of probabilistic uncertainty.
@@
Bayesian methods provide a natural framework for uncertainty quantification but have seen limited integration with deep survival models. GP-based approaches [10, 21] inherit GP limitations in scalability and high-dimensional processing. Recent works like BCESurv [26] explore bootstrap confidence intervals, yet these post-hoc approximations lack the coherence of Bayesian posteriors. Consequently, existing Bayesian survival models either sacrifice scalability for uncertainty quantification or compromise on model flexibility, leaving a critical gap in high-dimensional, uncertainty-aware survival analysis.
@@
CoxPH [7]
AFT [6]
RSF [19]
SSVM [36]
GP survival models [10, 21]
MTLR [43]
DeepHit [28]
DeepSurv [20]
Logistic Hazard [13]
CoxTime [25]
CoxCC [25]
PMF [24]
PCHazard [24]
BCESurv [26]
DySurv [32]
@@
Table A1: Summary of survival-analysis methods: uncertainty quantification, time domain, and deep-learning status.
@@
MTLR [43]
DeepHit [28]
DeepSurv [20]
Logistic Hazard [13]
CoxTime [25]
CoxCC [25]
PMF [24]
PCHazard [24]
BCESurv [26]
DySurv [32]
@@
MTLR [43]
DeepHit [28]
DeepSurv [20]
Logistic Hazard [13]
CoxTime [25]
CoxCC [25]
PMF [24]
PCHazard [24]
BCESurv [26]
DySurv [32]
@@
MTLR [43]
DeepHit [28]
DeepSurv [20]
Logistic Hazard [13]
CoxTime [25]
CoxCC [25]
PMF [24]
PCHazard [24]
BCESurv [26]
DySurv [32]
@@
MTLR [43]
DeepHit [28]
DeepSurv [20]
Logistic Hazard [13]
CoxTime [25]
CoxCC [25]
PMF [24]
PCHazard [24]
BCESurv [26]
DySurv [32]
@@
MTLR [43]
DeepHit [28]
DeepSurv [20]
Logistic Hazard [13]
CoxTime [25]
CoxCC [25]
PMF [24]
PCHazard [24]
BCESurv [26]
DySurv [32]
@@
CoxPH [7]
Weibull AFT [6]
RSF [19]
SSVM [36]
@@
CoxPH [7]
Weibull AFT [6]
RSF [19]
SSVM [36]
@@
CoxPH [7]
Weibull AFT [6]
RSF [19]
SSVM [36]
@@
Table A5: Performance comparison of traditional survival models over five different train/test splits of each dataset. $\uparrow$ indicates higher is better; $\downarrow$ indicates lower is better. The SSVM method does not provide estimates of the survival function; therefore, the predicted ranks are used for the corresponding C-index evaluations while the IPCW-IBS metric cannot be computed.

### v2

@@ Appendix I Computational Speed-Ups
The experiments were conducted on NVIDIA RTX A6000 GPUs with 48GB of memory.
Running time
Table A1 reports the running time for a single fold on the Colon dataset at varying sample sizes. All folds and datasets were processed in parallel across multiple GPUs to ensure consistent timing.
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
$N=25$
1L 2L-6U 2L-16U
0.104 0.102 0.100
0.164 0.139 0.140
0.082 0.098 0.072
0.077 0.079 0.076
0.139 0.167 0.114
0.130 0.118 0.094
0.129 0.119 0.094
0.083 0.102 0.083
0.081 0.082 0.092
0.050 0.064 0.040
0.058 0.058 0.060
0.019 0.020 0.024
0.621 1.165 0.566
$N=125$
1L 2L-6U 2L-16U
0.075 0.093 0.118
0.120 0.126 0.142
0.063 0.079 0.072
0.070 0.090 0.072
0.113 0.175 0.134
0.100 0.119 0.114
0.069 0.104 0.095
0.087 0.092 0.084
0.060 0.078 0.092
0.048 0.040 0.049
0.062 0.071 0.070
0.023 0.025 0.025
3.926 16.454 22.673
$N=250$
1L 2L-6U 2L-16U
0.075 0.099 0.091
0.122 0.145 0.135
0.071 0.080 0.101
0.072 0.078 0.079
0.136 0.135 0.186
0.107 0.132 0.149
0.072 0.087 0.125
0.091 0.091 0.091
0.062 0.087 0.079
0.048 0.045 0.041
0.073 0.073 0.078
0.029 0.030 0.030
7.373 87.098 141.389
Table A1: Inference runtime for the Colon dataset (in minutes). The central analysis presented in Table A5-A6 is for $N=125$ and a MNP with 2 layers (2L) and 16 units (16U). The ablation study presented in Table A7-A8 is for $N=250$ and a MNP with 2L and 16U.
@@
Traditional survival models often impose parametric or semi-parametric assumptions on the hazard function. The Accelerated Failure Time (AFT) model [Carroll2003] assumes a linear relationship between covariates and the logarithm of survival time, with parametric baseline distributions (e.g., Weibull). While interpretable, such models struggle with high-dimensional data and nonlinear covariate effects. The Cox Proportional Hazards (CoxPH) model [Cox1972], a semi-parametric approach, avoids specifying the baseline hazard but assumes proportional hazards. Though widely adopted, CoxPH’s linear predictor and proportionality constraints limit its flexibility in complex data regimes.
@@
To mitigate parametric assumptions, non-parametric methods like Random Survival Forests (RSF) [Ishwaran2008] and Survival Support Vector Machines (SSVM) [Polsterl2015] emerged. RSF leverages ensemble learning for risk stratification but faces challenges in high-dimensional settings due to greedy tree induction. GP survival models [teh_survival] offer flexibility by modeling the hazard function nonparametrically, with inherent uncertainty quantification. Existing work has sought to address the cubic complexity in sample size of GPs by introducing variational inference techniques [Kim2018]. However, GPs remain fundamentally limited in scalability, particularly struggling with high-dimensional inputs and lacking the capacity to learn hierarchical representations, such as those required in image-based tasks [Salimbeni2017].
@@
The advent of deep learning revolutionized survival analysis by enabling automatic feature learning from high-dimensional inputs. DeepSurv [Katzman2018] extended CoxPH with neural networks, while DeepHit [Lee2018] employed multi-task learning for competing risks via discrete-time hazards. Discrete-time methods, including MTLR [Yu2011] and PCHazard [Kvamme2021], discretize the time axis to simplify likelihood computation, with recent advances like DySurv [Mesinovic2024] incorporating conditional variational inference for dynamic prediction. Sumo‑Net [rindt2022] introduces a partially monotonic NN that directly optimizes the right‑censored log‑likelihood, which is proven to be a strictly proper scoring rule—achieving strong log‑likelihood performance. DQS [Yanagisaw2023] formulates survival prediction using extensions of strictly proper scoring rules that remain proper under discrete-time survival settings. Despite their predictive prowess, these models rely on frequentist training, yielding point estimates without uncertainty quantification, a significant shortcoming in safety-critical applications. Comprehensive reviews [Wiegrebe2024] highlight the rapid growth of deep survival methods but underscore their neglect of probabilistic uncertainty.
@@
Bayesian methods provide a natural framework for uncertainty quantification but have seen limited integration with deep survival models. GP-based approaches [teh_survival, Kim2018] inherit GP limitations in scalability and high-dimensional processing. Recent works like BCESurv [Kvamme2023] explore bootstrap confidence intervals, yet these post-hoc approximations lack the coherence of Bayesian posteriors. Consequently, existing Bayesian survival models either sacrifice scalability for uncertainty quantification or compromise on model flexibility, leaving a critical gap in high-dimensional, uncertainty-aware survival analysis.
@@
CoxPH [Cox1972]
AFT [Carroll2003]
RSF [Ishwaran2008]
SSVM [Polsterl2015]
GP survival models [teh_survival, Kim2018]
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo‑Net [rindt2022]
DQS [Yanagisaw2023]
@@
Table A2: Summary of Survival Analysis methods: Bayesian Uncertainty Quantification, Time Domain, and Deep‐Learning Status.
@@
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
@@
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
$N=25$
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.032
0.000 ($\times$) 0.029
0.000 ($\times$) 0.009
0.000 ($\times$) 0.037
0.000 ($\times$) 0.010
0.000 ($\times$) 0.008
0.000 ($\times$) 0.032
0.000 ($\times$) 0.037
0.000 ($\times$) 0.022
0.000 ($\times$) 0.201
0.132 ($\checkmark$) 0.011
0.000 ($\times$) 0.051
0.833 ($\checkmark$) 0.014
$N=50$
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.056
0.000 ($\times$) 0.168
0.000 ($\times$) 0.036
0.000 ($\times$) 0.062
0.000 ($\times$) 0.021
0.000 ($\times$) 0.028
0.000 ($\times$) 0.071
0.000 ($\times$) 0.073
0.000 ($\times$) 0.014
0.000 ($\times$) 0.241
0.051 ($\checkmark$) 0.036
0.000 ($\times$) 0.048
0.539 ($\checkmark$) 0.012
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
$N=100$
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.034
0.000 ($\times$) 0.030
0.000 ($\times$) 0.003
0.000 ($\times$) 0.034
0.000 ($\times$) 0.011
0.000 ($\times$) 0.001
0.000 ($\times$) 0.033
0.000 ($\times$) 0.056
0.000 ($\times$) 0.038
0.000 ($\times$) 0.172
0.683 ($\checkmark$) 0.004
0.005 ($\times$) 0.032
0.419 ($\checkmark$) 0.004
$N=150$
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.035
0.000 ($\times$) 0.040
0.007 ($\times$) 0.003
0.000 ($\times$) 0.062
0.346 ($\checkmark$) 0.005
0.001 ($\times$) 0.001
0.000 ($\times$) 0.034
0.000 ($\times$) 0.066
0.000 ($\times$) 0.034
0.000 ($\times$) 0.183
0.345 ($\checkmark$) 0.003
0.000 ($\times$) 0.031
0.639 ($\checkmark$) 0.004
Table A4: Performance comparison of survival models over synthetic data (part 2). A checkmark ($\checkmark$) indicates that the null hypothesis of perfect D-Calibration was not rejected at $\alpha=0.05$ (model considered well-calibrated); a cross ($\times$) indicates rejection of D-Calibration (model considered not well-calibrated). The best results for the KM-Calibration are shown in bold, and the second-best results are underlined. $\downarrow$ indicates lower is better.
M.2 Real Data Experiment
M.2.1 Central Analysis
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
@@
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
@@
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
@@
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
COLON
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.016
0.001 ($\times$) 0.089
0.047 ($\times$) 0.024
0.002 ($\times$) 0.019
0.014 ($\times$) 0.011
0.011 ($\times$) 0.011
0.001 ($\times$) 0.017
0.007 ($\times$) 0.029
0.000 ($\times$) 0.012
0.000 ($\times$) 0.362
0.741 ($\checkmark$) 0.014
0.381 ($\checkmark$) 0.017
0.594 ($\checkmark$) 0.020
METABRIC
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.023
0.101 ($\checkmark$) 0.064
0.006 ($\times$) 0.012
0.000 ($\times$) 0.026
0.012 ($\times$) 0.012
0.013 ($\times$) 0.019
0.002 ($\times$) 0.017
0.008 ($\times$) 0.036
0.000 ($\times$) 0.021
0.000 ($\times$) 0.306
0.600 ($\checkmark$) 0.008
0.135 ($\checkmark$) 0.020
0.661 ($\checkmark$) 0.012
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
GBSG
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.011
0.000 ($\times$) 0.149
0.084 ($\checkmark$) 0.009
0.008 ($\times$) 0.017
0.247 ($\checkmark$) 0.009
0.256 ($\checkmark$) 0.005
0.003 ($\times$) 0.015
0.063 ($\checkmark$) 0.024
0.001 ($\times$) 0.015
0.000 ($\times$) 0.369
0.524 ($\checkmark$) 0.010
0.234 ($\checkmark$) 0.012
0.735 ($\checkmark$) 0.009
NWTCO
D-Calibration (p-value) KM-Calibration $\downarrow$
0.570 ($\checkmark$) 0.011
0.575 ($\checkmark$) 0.014
0.887 ($\checkmark$) 0.003
0.397 ($\checkmark$) 0.012
0.883 ($\checkmark$) 0.003
0.954 ($\checkmark$) 0.003
0.487 ($\checkmark$) 0.013
0.697 ($\checkmark$) 0.008
0.312 ($\checkmark$) 0.010
0.000 ($\times$) 0.260
0.991 ($\checkmark$) 0.004
0.916 ($\checkmark$) 0.005
0.920 ($\checkmark$) 0.007
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
WHAS
D-Calibration (p-value) KM-Calibration $\downarrow$
0.001 ($\times$) 0.035
0.195 ($\checkmark$) 0.067
0.076 ($\checkmark$) 0.025
0.013 ($\times$) 0.033
0.035 ($\times$) 0.017
0.106 ($\checkmark$) 0.019
0.003 ($\times$) 0.031
0.125 ($\checkmark$) 0.022
0.000 ($\times$) 0.025
0.000 ($\times$) 0.281
0.735 ($\checkmark$) 0.021
0.081 ($\checkmark$) 0.033
0.335 ($\checkmark$) 0.031
SUPPORT
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.099
0.000 ($\times$) 0.124
0.000 ($\times$) 0.015
0.000 ($\times$) 0.095
0.000 ($\times$) 0.016
0.000 ($\times$) 0.020
0.000 ($\times$) 0.084
0.000 ($\times$) 0.070
0.000 ($\times$) 0.088
0.000 ($\times$) 0.099
0.143 ($\checkmark$) 0.017
0.002 ($\times$) 0.038
0.063 ($\checkmark$) 0.083
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
VLC
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.072
0.000 ($\times$) 0.107
0.004 ($\times$) 0.006
0.000 ($\times$) 0.078
0.021 ($\times$) 0.010
0.004 ($\times$) 0.011
0.000 ($\times$) 0.077
0.000 ($\times$) 0.080
0.000 ($\times$) 0.074
0.077 ($\checkmark$) 0.028
0.131 ($\checkmark$) 0.011
0.000 ($\times$) 0.054
0.436 ($\checkmark$) 0.013
SAC3
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.020
0.003 ($\times$) 0.094
0.000 ($\times$) 0.020
0.000 ($\times$) 0.039
0.000 ($\times$) 0.014
0.000 ($\times$) 0.016
0.000 ($\times$) 0.016
0.000 ($\times$) 0.029
0.000 ($\times$) 0.021
0.000 ($\times$) 0.146
0.478 ($\checkmark$) 0.016
0.021 ($\times$) 0.034
0.624 ($\checkmark$) 0.016
Table A6: Performance comparison of deep survival models over five different train/test splits of each dataset (part 2). A checkmark ($\checkmark$) indicates that the null hypothesis of perfect D-Calibration was not rejected at $\alpha=0.05$ (model considered well-calibrated); a cross ($\times$) indicates rejection of D-Calibration (model considered not well-calibrated). The best results for the KM-Calibration are shown in bold, and the second-best results are underlined. $\downarrow$ indicates lower is better.
M.2.2 Ablation Study with $N=250$
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
@@
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
COLON
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.007
0.000 ($\times$) 0.068
0.001 ($\times$) 0.006
0.000 ($\times$) 0.012
0.022 ($\times$) 0.005
0.001 ($\times$) 0.003
0.000 ($\times$) 0.008
0.001 ($\times$) 0.018
0.000 ($\times$) 0.012
0.000 ($\times$) 0.376
0.323 ($\checkmark$) 0.007
0.086 ($\checkmark$) 0.009
0.404 ($\checkmark$) 0.011
METABRIC
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.017
0.044 ($\times$) 0.072
0.001 ($\times$) 0.007
0.000 ($\times$) 0.017
0.001 ($\times$) 0.008
0.000 ($\times$) 0.006
0.000 ($\times$) 0.008
0.003 ($\times$) 0.027
0.000 ($\times$) 0.013
0.000 ($\times$) 0.273
0.192 ($\checkmark$) 0.007
0.137 ($\checkmark$) 0.013
0.708 ($\checkmark$) 0.013
method
Method
MTLR [Yu2011]
DeepHit [Lee2018]
DeepSurv [Katzman2018]
Logistic Hazard [Gensheimer2019]
CoxTime [Kvamme2019]
CoxCC [Kvamme2019]
PMF [Kvamme2021]
PCHazard [Kvamme2021]
BCESurv [Kvamme2023]
DySurv [Mesinovic2024]
Sumo-Net  [rindt2022]
DQS [Yanagisaw2023]
NeuralSurv (Ours)
GBSG
D-Calibration (p-value) KM-Calibration $\downarrow$
0.000 ($\times$) 0.010
0.159 ($\checkmark$) 0.134
0.052 ($\checkmark$) 0.006
0.000 ($\times$) 0.016
0.171 ($\checkmark$) 0.004
0.059 ($\checkmark$) 0.006
0.000 ($\times$) 0.010
0.004 ($\times$) 0.015
0.000 ($\times$) 0.013
0.000 ($\times$) 0.336
0.337 ($\checkmark$) 0.004
0.084 ($\checkmark$) 0.008
0.617 ($\checkmark$) 0.003
Table A8: Performance comparison of deep survival models on the ablation study with 250 observations, over five different train/test splits of each dataset (part 2). A checkmark ($\checkmark$) indicates that the null hypothesis of perfect D-Calibration was not rejected at $\alpha=0.05$ (model considered well-calibrated); a cross ($\times$) indicates rejection of D-Calibration (model considered not well-calibrated). The best results for the KM-Calibration are shown in bold, and the second-best results are underlined. $\downarrow$ indicates lower is better.
M.2.3 Comparison to Traditional Survival Models
method
Method
CoxPH [Cox1972]
Weibull AFT [Carroll2003]
RSF [Ishwaran2008]
SSVM [Polsterl2015]
@@
CoxPH [Cox1972]
Weibull AFT [Carroll2003]
RSF [Ishwaran2008]
SSVM [Polsterl2015]
@@
CoxPH [Cox1972]
Weibull AFT [Carroll2003]
RSF [Ishwaran2008]
SSVM [Polsterl2015]
@@
Table A9: Performance comparison of traditional survival models over five different train/test splits of each dataset. $\uparrow$ indicates higher is better; $\downarrow$ indicates lower is better. The SSVM method does not provide estimates of the survival function; the predicted ranks are used for the corresponding C-index evaluations while the IPCW-IBS metric cannot be computed.
method
Method
CoxPH [Cox1972]
Weibull AFT [Carroll2003]
RSF [Ishwaran2008]
COLON
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.913 ($\checkmark$) 0.006
0.788 ($\checkmark$) 0.010
0.791 ($\checkmark$) 0.009
NWTCO
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.979 ($\checkmark$) 0.003
0.986 ($\checkmark$) 0.004
0.999 ($\checkmark$) 0.002
GBSG
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.950 ($\checkmark$) 0.004
0.767 ($\checkmark$) 0.008
0.854 ($\checkmark$) 0.003
method
Method
CoxPH [Cox1972]
Weibull AFT [Carroll2003]
RSF [Ishwaran2008]
METABRIC
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.846 ($\checkmark$) 0.008
0.759 ($\checkmark$) 0.006
0.746 ($\checkmark$) 0.009
WHAS
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.730 ($\checkmark$) 0.017
0.650 ($\checkmark$) 0.018
0.625 ($\checkmark$) 0.022
SUPPORT
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.354 ($\checkmark$) 0.010
0.420 ($\checkmark$) 0.012
0.593 ($\checkmark$) 0.010
method
Method
CoxPH [Cox1972]
Weibull AFT [Carroll2003]
RSF [Ishwaran2008]
VLC
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.597 ($\checkmark$) 0.008
0.759 ($\times$) 0.007
0.414 ($\times$) 0.013
SAC3
D-Calibration KM-Calibration $\downarrow$
(p-value)
0.357 ($\checkmark$) 0.005
0.038 ($\checkmark$) 0.018
0.706 ($\checkmark$) 0.012
Table A10: Performance comparison of traditional survival models over five different train/test splits of each dataset (part 2). A checkmark ($\checkmark$) indicates that the null hypothesis of perfect D-Calibration was not rejected at $\alpha=0.05$ (model considered well-calibrated); a cross ($\times$) indicates rejection of D-Calibration (model considered not well-calibrated). The best results for the KM-Calibration are shown in bold, and the second-best results are underlined. $\downarrow$ indicates lower is better.
M.3 Prior Sensitivity Analysis
method
Gamma Prior of $\phi$
$(\alpha_{0}=1,\beta_{0}=1)$
$(\alpha_{0}=2,\beta_{0}=2)$
$(\alpha_{0}=0.5,\beta_{0}=0.5)$
COLON
Gamma Posterior of $\phi$ Posterior Median C-index $\uparrow$ IPCW IBS $\downarrow$ D-Calibration KM-Calibration $\downarrow$
and 95% CI (p-value)
($\tilde{\alpha}$ = 60.877, $\tilde{\beta}$ = 78.789) 0.768 [0.591, 0.978] 0.671 0.218 0.594 ($\checkmark$) 0.020
($\tilde{\alpha}$ = 45.865, $\tilde{\beta}$ = 79.789) 0.571 [0.421, 0.753] 0.593 0.237 0.601 ($\checkmark$) 0.025
($\tilde{\alpha}$ = 38.333, $\tilde{\beta}$ = 78.289) 0.485 [0.347, 0.656] 0.512 0.229 0.715 ($\checkmark$) 0.023
method
Gamma Prior of $\phi$
$(\alpha_{0}=1,\beta_{0}=1)$
$(\alpha_{0}=2,\beta_{0}=2)$
$(\alpha_{0}=0.5,\beta_{0}=0.5)$
METABRIC
Gamma Posterior of $\phi$ Posterior Median C-index $\uparrow$ IPCW IBS $\downarrow$ D-Calibration KM-Calibration $\downarrow$
and 95% CI (p-value)
($\tilde{\alpha}$ = 48.406, $\tilde{\beta}$ = 70.176) 0.685 [0.509, 0.897] 0.584 0.212 0.661 ($\checkmark$) 0.012
($\tilde{\alpha}$ = 48.034, $\tilde{\beta}$ = 71.176) 0.670 [0.498, 0.879] 0.536 0.201 0.819 ($\checkmark$) 0.009
($\tilde{\alpha}$ = 47.347, $\tilde{\beta}$ = 69.676) 0.675 [0.500, 0.886] 0.553 0.200 0.802 ($\checkmark$) 0.009
method
Gamma Prior of $\phi$
$(\alpha_{0}=1,\beta_{0}=1)$
$(\alpha_{0}=2,\beta_{0}=2)$
$(\alpha_{0}=0.5,\beta_{0}=0.5)$
GBSG
Gamma Posterior of $\phi$ Posterior Median C-index $\uparrow$ IPCW IBS $\downarrow$ D-Calibration KM-Calibration $\downarrow$
and 95% CI (p-value)
($\tilde{\alpha}$ = 58.650, $\tilde{\beta}$ = 85.462) 0.682 [0.522, 0.873] 0.657 0.188 0.735 ($\checkmark$) 0.009
($\tilde{\alpha}$ = 62.386, $\tilde{\beta}$ = 86.462) 0.718 [0.554, 0.911] 0.602 0.195 0.808 ($\checkmark$) 0.010
($\tilde{\alpha}$ = 57.549, $\tilde{\beta}$ = 84.962) 0.673 [0.514, 0.863] 0.665 0.189 0.772 ($\checkmark$) 0.010
Table A11: Prior sensitivity analysis on $\phi$ using priors with double and half the original variance..